OpenStudy (gabylovesyou):

After a discount of 10%, the price of a pair of jeans, including tax, was $69.62. The following equation can be used to find x, the price of the pair of jeans without the discount. x - 0.1x + 0.07(x - 0.1x) = 69.62 What is the price of the pair of jeans without the discount? $70.58 $72.29 $74.86 $80.02

OpenStudy (anonymous):

$72.29 I'm going to show you all the steps: x - 0.1x + 0.07(x - 0.1x) = 69.62 x - 0.1x + 0.07x - 0.007x = 69.62 x - 0.03x - 0.007x = 69.62 x - 0.037x = 69.62 .963x = 69.62 x = 69.62/.963 x= 72.29